What is the probability of getting a 3 or odd number on a number cube ?
Answer by stanbon(75887)

What is the probability of getting a 3 or odd number on a number cube ?
----
P(3 or odd) = P(3)+P(odd)-P(3 and odd)
= 1/6 + 3/6 - 1/6
= 1/6
======================
Cheers,
Stan H.